# TCIP: Threshold-Controlled Iterative Pyramid Network for Deformable Medical Image Registration

### v1

@@ 3.1 Overview of TCIP
3.2 Feature-Enhanced Residual Module (FERM)
To effectively capture anatomical structure details while learning to suppress irrelevant features, we propose the Feature-Enhanced Residual Module (FERM) as the core component of the decoder for estimating deformation fields. As shown in Figure 3, FERM comprises three blocks, namely Feature Fusion Block (FFB), Squeeze Excitation Block (SEB), and Deformation Field Estimator (DeF). Specifically, for the fourth decoding layer (leftmost decoding layer in Figure 2), FFB directly fuses $F_{4}$ and $M_{4}$, and then extracts the fused features $R_{4}$. For layers $l\in\{1,2,3\}$, FFB first upsamples the deformation field $\phi_{l+1}$ by a factor of 2 using linear interpolation to obtain $\hat{\phi}_{l+1}$ at the scale of $M_{l}$, and then warps $M_{l}$ to obtain $M_{l}^{w}$. Then, FFB fuses $F_{l}$ and $M_{l}^{w}$ and extracts the features $R_{l}\in\mathbb{R}^{C\times H\times W\times D}$, where $C,H,W,D$ denote the number of channels, height, width, and depth of the image, respectively. Formally,
$R_{l}=\left\{\begin{array}[]{ll}\text{FFB}\left(|M_{4},F_{4}|\right),&l=4,\\ \text{FFB}\left(|M_{l}^{w},F_{l}|\right),&l\in\{1,2,3\},\end{array}\right$. (1)
where $|\cdot|$ denotes the concatenation operator. FFB comprises two 3D convolutional layers $\mathcal{C}1$ and $\mathcal{C}2$, each followed by a LeakyReLU activation function $\gamma$. In the first convolutional layer $\mathcal{C}1$, FFB reduces the concatenated channel dimension from $2C$ to $C$. A residual connection adds the output of the first activation to the second convolutional layer output, preserving original image features while enabling the extraction of high‑level anatomical structure details. The FFB function is formally defined as follows:
$\text{FFB($\cdot$)}=\gamma(\mathcal{C}2_{C\rightarrow C}(\gamma(\mathcal{C}1_{2C\rightarrow C}(\cdot)))+\gamma(\mathcal{C}1_{2C\rightarrow C}(\cdot))).$ (2)
Subsequently, FERM extends the 2D SEB (Hu, Shen, and Sun 2018) to a 3D variant to suppress irrelevant features in $R_{l}$, thereby mitigating propagation and accumulation of anatomical misalignment. Specifically, SEB first employs global average pooling to compress spatial dimensions and extract global informative features, producing the channel descriptor vector $Z_{l}\in\mathbb{R}^{C}$ as follows:
where $c\in\{1,\cdots,C\}$ Then, SEB applies two fully connected layers with LeakyReLU $\gamma$ and sigmoid $\sigma$ activation functions to compute the channel weight vector $S_{l}$ as follows:
$S_{l}=\sigma({W}_{2}\gamma({W}_{1}Z_{l})),$ (4)
where ${W}_{1}\in\mathbb{R}^{\frac{C}{r}\times C}$ and ${W}_{2}\in\mathbb{R}^{C\times\frac{C}{r}}$. The hyperparameter $r$ denotes the reduction ratio, which is used to gauge the computational cost. The first linear layer reduces the dimensionality of $Z_{l}$ to $\frac{C}{r}$, for efficient channel-wise compression and information aggregation. The second linear layer then restores the dimensionality back to $C$. By adopting this design, SEB extracts the global contextual information from each channel through a compression process. SEB then models inter-channel dependencies and recalibrates the channels via an expansion process, thereby capturing global dependencies between channels. Subsequently, the resulting weight vector $S_{l}$ is applied to $R_{l}$ via per‑channel element‑wise multiplication $\ast$, yielding the enhanced feature map $O_{l}\in\mathbb{R}^{C\times H\times W\times D}$, defined as follows:
$O_{l,c}=R_{l,c}\ast S_{l,c},c\in\{1,\cdots,C\},$ (5)
By differentiating channels with associated weights (i.e., $S_{l}$), SEB emphasizes deformation‑relevant features while suppressing irrelevant information (see Table 5).
Finally, DeF processes $O_{l}$ via two convolutional layers $\mathcal{C}3$ and $\mathcal{C}4$ to estimate the deformation field $\phi_{l}$ as follows:
$\phi_{l}=\mathrm{\mathcal{C}4_{\frac{C}{2}\rightarrow 3}}(\mathrm{\mathcal{C}3_{C\rightarrow\frac{C}{2}}}(O_{l})).$ (6)
Layer $\mathcal{C}3$ halves the channel dimension from $C$ to $\frac{C}{2}$, while $\mathcal{C}4$ outputs three channels corresponding to the 3D deformation field. Notably, unlike the decoder adopted by IIRP (Ma et al. 2024), which relies solely on multiple sequential residual blocks to estimate the deformation field, FERM integrates lightweight channel attention alongside a minimal number of residual blocks, designed to yield elevated performance with fewer parameters (see Table 1).
3.3 Threshold-Controlled Iterative (TCI)
@@
The number of iterations is crucial for optimizing the deformation field; early termination may yield suboptimal accuracy, while excessive iterations can introduce misalignment and consume unnecessary computing resources. To adaptively determine the number of iterations for varying image pairs, we propose a dual-stage TCI strategy to evaluate both stability and convergence using historical registration results to dynamically determine when to stop. As shown in Figure 4, at the $l$th decoding layer, TCI employs a sliding window $W_{l}$ to dynamically track the similarity differences of $t$ number of recently estimated registration results.
@@
1:Input: $(l+1)$th deformation field $\hat{\phi}_{l+1}$, feature maps $F_{l}$, $M_{l}$, original images $I_{f}$, $I_{m}$, $W_{l}$, $k\leftarrow 1$
2:Output: $l$th deformation field $\phi_{l}$
@@
5:  $M_{l}^{w}\leftarrow M_{l}\circ\varphi_{l}^{k-1}$ #If $l=4$, this step is skipped
6:  $\varphi_{l}^{k}\leftarrow\text{FERM}(M_{l}^{w},F_{l})$ #$l=4$, replace $M_{l}^{w}$ with $M_{4}$
7:  $\varphi_{l}^{k}\leftarrow{\varphi}_{l}^{k-1}\circ\varphi_{l}^{k}+\varphi_{l}^{k}$ #If l=4&k=1l=4\&k=1, this step is skipped. If $l\neq 4\&k=1$, replace ${\varphi}_{l}^{k-1}$ with $\hat{\phi}_{l+1}$
8:  $\hat{\varphi}_{l}^{k}\leftarrow\text{upsample}(\varphi_{l}^{k},2^{l-1})$
9:  $s_{l}^{k}\leftarrow\text{sim}(I_{f},I_{m}\circ\hat{\varphi}_{l}^{k})$ #see (7)
10:  if $\text{len}(W_{l})\geq t$ then#TCI
11:   $\varepsilon_{l}\leftarrow\text{std}\left(W_{l}\right)$ #see (8)
12:   if $\varepsilon_{l}\leq\delta_{s}$ then #Stage 1: assess stability
13:    $\Delta s\leftarrow s_{l}^{k}-s_{l}^{k-1}$ #see (9)
14:    if $\Delta s\leq\delta_{c}$ then #Stage 2: assess convergence
15:     return $\hat{\phi}_{l}\leftarrow\text{upsample}(\varphi_{l}^{k},2)$
16:  append $s_{l}^{k}$ to $W_{l}$ #If $\text{len}(W_{l})=t+1$, remove $s_{l}^{k-t}$ from $W_{l}$
17:  $k\leftarrow k+1$
18:end while
To compute the similarity difference $s_{l}^{k}$ between the warped image $I_{m}^{w}$ and the fixed image $I_{f}$ at the $k$th iteration, TCI first upsamples the deformation field $\varphi_{l}^{k}$ estimated by DeF, yielding the upsampled deformation field $\widehat{\varphi}_{l}^{k}$ at the original moving image resolution. The upsampling is implemented via linear interpolation with a factor of $2^{l-1}$. Next, the upsampled deformation field $\widehat{\varphi}_{l}^{k}$ provides a displacement vector for each voxel and is applied to the moving image $I_{m}$ to produce the corresponding warped image $I_{m}^{w}$, i.e., $I_{m}^{w}=I_{m}\circ\widehat{{\varphi}}_{l}^{k}$, where $\circ$ denotes the deformation operation, implemented via the Spatial Transformer Network (Jaderberg et al. 2015). Finally, $s_{l}^{k}$ is computed as follows:
$s_{l}^{k}=\text{sim}\left(I_{f},I_{m}^{w}\right),$ (7)
where $\mathrm{sim}(\cdot)$ denotes the similarity metric. In this work, we always adopt Normalized Cross-Correlation (NCC) to compute the registration similarity after examining various candidates (see Appendix F).
After obtaining the similarity differences of $t$ number of recently estimated registration results, TCI evaluates their stability and convergence in a dual-stage process, respectively. In the first stage, TCI assesses stability by computing the standard deviation $\varepsilon_{l}$ of similarity differences $s_{l}^{k}$ within the window $W_{l}=\{s_{l}^{k-t+1},\cdots,s_{l}^{k}\}$, defined as follows:
$\varepsilon_{l}=\text{std}\left(W_{l}\right).$ (8)
If $\varepsilon_{l}$ falls below the predefined stability threshold $\delta_{s}$, the most recent $t$ number of deformation fields are deemed stable. TCI then proceeds to the subsequent convergence evaluation stage. Otherwise, an additional iteration is performed without the necessity to continue the following stage.
@@
If $\Delta s$ falls below the predefined convergence threshold $\delta_{c}$, the deformation field at the $k$th iteration is deemed converged, the iteration terminates, and the deformation field $\phi_{l}$ is passed to the (l1)(l-1)th layer as input (see Figure 2). Otherwise, additional iterations continue at the $l$th layer. Following (Ma et al. 2024), TIC also leverages a predefined maximum number of iterations $k_{max}$ to avoid rare non-convergence situations. The pseudocode of TCI is presented in Algorithm 1.
Notably, our TCI fundamentally differs from IIRP’s single‑stage method (Ma et al. 2024), which performs only one similarity check between two consecutive registrations. Such a single-stage check, constrained by the stochastic nature of convergence, may lead to premature termination. In contrast, TCI introduces a two-stage evaluation strategy. Owing to the additional stability check, TCI makes more informed iteration‑stopping decisions while maintaining comparable inference time (see Table 4). In addition, because IIRP’s decoder comprises multiple residual blocks, applying a single‑stage strategy during training is highly memory‑intensive; consequently, IIRP activates this strategy during inference only. By contrast, TCIP’s lightweight FERM decoding layer enables the integration of TCI for both training and inference. This joint optimization of the network and iteration controller from the commencement of training enhances registration performance without prolonging the computation significantly (see Tables 1 and  2).
3.4 Loss Function
Being an unsupervised registration network, similar to the prior SOTA model (Ma et al. 2024), TCIP employs NCC to evaluate the loss between the fixed image $I_{f}$ and the warped moving image $I_{w}^{m}$, which is defined as follows:
$\mathcal{L}_{ncc}(I_{f},I_{m}^{w})=$ (10)
$-\sum_{p\in\Omega}\frac{\sum_{p_{i}}(I_{f}(p_{i})-\overline{I_{f}}(p))(I_{m}^{w}(p_{i})-\overline{I_{m}^{w}}(p))}{\sqrt{\sum_{p_{i}}(I_{f}(p_{i})-\overline{I_{f}}(p))^{2}\sum_{p_{i}}(I_{m}^{w}(p_{i})-\overline{I_{m}^{w}}(p))^{2}}},$
@@ 3.4 Loss Function
Dice(%)↑ Time(s) $|{J}_{\phi}|_{<0}$ Dice(%)↑ Time(s) $|{J}_{\phi}|_{<0}$ Dice(%)↑ Time(s) $|{J}_{\phi}|_{<0}$
U-Net
VM (CVPR’18) 56.0 0.09 <<0.7% 65.1 0.02 <<0.3% 75.9 0.01 <<0.2% 319.4K
@@
TCIP (ours) 66.4 0.16 <<0.004% 73.3 0.13 <<0.001% 80.5 0.09 <<0.004% 330.4K
Moreover, to enforce the smoothness of the deformation field, we incorporate the $L_{2}$-norm $||\cdot||$ on the gradient of $\phi$:
$\mathcal{L}_{smooth}(\phi)=\sum\nolimits_{p\in\Omega}\lVert\nabla\phi(p)\rVert^{2}.$ (11)
Therefore, the total loss in TCIP is defined as follows:
$\mathcal{L}_{total}(I_{f},I_{m})=\mathcal{L}_{ncc}(I_{f},I_{m}^{w})+\lambda\mathcal{L}_{smooth}(\phi),$ (12)
where $\lambda$ denotes the hyperparameter used to leverage the loss function.
@@ 4 Experimental Results
To assess the performance of TCIP, we conduct extensive experiments on three public MRI datasets, namely LPBA (Shattuck et al. 2008), IXI (Avants et al. 2009), and Mindboggle (Klein and Tourville 2012), and one abdomen CT dataset FLARE (Ma et al. 2022). Detailed descriptions of these datasets, evaluation metrics, and experimental setups, are provided in Appendix A.
4.1 Comparison with Baseline Models
Figure 5: Visualization of registration results and corresponding deformation fields across different models.
To benchmark the performance of TCIP, we choose two U-Net models, namely VM (Balakrishnan et al. 2019) and TM (Chen et al. 2022), two cascaded models, namely RCN (Zhao et al. 2019a) and SDH (Zhou et al. 2023), and six multi-scale pyramid models, namely RDN (Hu et al. 2022), PR++ (Kang et al. 2022), I2G (Liu et al. 2022), RDP (Wang, Ni, and Wang 2024), IIRP (Ma et al. 2024), and SACB (Cheng et al. 2025) as baseline models. Table 1 presents the performance of different registration models on the Mindboggle, LPBA, and IXI datasets, measured using the Dice (Dice 1945), Jacobian determinant ($|{J}_{\phi}|_{<0}$), and inference time (s) metrics. Appendix B presents HD95 (Huttenlocher, Klanderman, and Rucklidge 1993), ASSD (Taha and Hanbury 2015), and MSE metrics for these registration models across all three datasets. Table 2 presents the performance of different registration models (competing pyramid models in Table 1) on the FLARE dataset. The experimental results shown in both tables demonstrate that TCIP outperforms all the other registration models on the Dice metric across all four datasets. Unlike SACB, which estimates the deformation field only once per decoding layer, TCIP employs TCI to adaptively determine the number of iterations at each layer, thereby yielding superior performance, achieving Dice improvements of 1.3%, 0.2%, 0.2%, and 0.5% on the four datasets, respectively. Notably, RDP and IIRP also employ iterative strategies. However, RDP adopts a fixed number of iterations and IIRP relies on a single-stage stopping criterion evaluation, both of which lead to inferior performance. In contrast, our TCIP employs TCI with a dual-stage strategy at each decoding layer, thereby achieving elevated performance while incurring only a marginal increase in inference time over IIRP. Furthermore, FERM is designed to be efficient to excel at extracting anatomical structure details and suppresses irrelevant features, thereby enabling TCIP to achieve superior performance with only approximately 80% of the parameters needed by IIRP.
@@ 4.1 Comparison with Baseline Models
Model Dice (%)↑ Time (s) $|{J}_{\phi}|_{<0}$
@@
TCIP (ours) 74.8 0.06 <<0.2%
Figure 5 visualizes the registration results on the Mindboggle and LPBA datasets. Performance differences among models are more noticeable in the three structural regions of the brain images, highlighted by red, yellow, and blue boxes. These visualizations further demonstrate that the moving images, after being warped by the deformation fields produced by TCIP, align more closely with fixed images. Additionally, we perform quantitative analyses of specific regions in the left and right hemispheres to further illustrate variations in anatomical structure registration performance across different models. The results, as presented in Appendix C, further demonstrate the advantages of TCIP.
4.2 Generalizability of FERM and TCI
Table 3: The generalizability of FERM on the IXI dataset
Model Dice (%) MSE ($10^{-3}$)
I2G (+ FERM) 79.3 (79.7) 0.40 (0.39)
PR++ (+ FERM) 79.3 (79.4) 0.36 (0.37)
IIRP (+ FERM) 80.4 (81.2) 0.23 (0.19)
To demonstrate the generalizability of FERM and TCI, we integrate each into existing registration models and report the results in Tables 3 and 4, respectively. Table 3 shows that replacing the decoder with FERM yields consistent performance gains across all three pyramid models, showcasing FERM’s superior ability to suppress irrelevant features. Furthermore, Table 4 demonstrates that substituting IIRP’s original iterative strategy with TCI yields performance improvement, underscoring the advantage of TCI in determining the number of iteration dynamically.

### v2

@@ 3.2 Threshold-Controlled Iterative (TCI) Strategy
The number of iterations is crucial for optimizing the deformation field; early termination may yield suboptimal accuracy, while excessive iterations may introduce misalignment and consume unnecessary computing resources. To adaptively determine the number of iterations for varying image pairs, we propose a dual-stage TCI strategy to evaluate both stability and convergence using historical registration results to dynamically determine when to stop. As shown in Figure 4, at the $l$th decoding layer, TCI employs a sliding window $W_{l}$ to dynamically track the similarity differences of $t$ number of recently estimated registration results.
To compute the similarity difference $s_{l}^{k}$ between the warped image $I_{m}^{w}$ and the fixed image $I_{f}$ at the $k$th iteration, TCI first upsamples the deformation field $\varphi_{l}^{k}$, yielding the upsampled deformation field $\tilde{\varphi}_{l}^{k}$ at the original moving image resolution. The upsampling is implemented via linear interpolation with a factor of $2^{l-1}$. Next, $\tilde{\varphi}_{l}^{k}$ provides a displacement vector for each voxel to be applied to the moving image $I_{m}$ to produce the corresponding warped image $I_{m}^{w}$, i.e., $I_{m}^{w}=I_{m}\circ\tilde{\varphi}_{l}^{k}$. Formally,
$s_{l}^{k}=\text{sim}\left(I_{f},I_{m}^{w}\right),$ (14)
where $\mathrm{sim}(\cdot)$ denotes the similarity metric. In this work, we always adopt Normalized Cross-Correlation (NCC) to compute the registration similarity after examining various candidates (see Appendix I).
After obtaining the similarity differences of $t$ number of recently estimated registration results, TCI evaluates their stability and convergence in a dual-stage process, respectively. In the first stage, TCI assesses stability by computing the standard deviation $\varepsilon_{l}$ of similarity differences $s_{l}^{k}$ within the window $W_{l}=\{s_{l}^{k-t+1},\cdots,s_{l}^{k}\}$ as follows:
$\varepsilon_{l}=\text{std}\left(W_{l}\right).$ (15)
If $\varepsilon_{l}$ falls below the predefined stability threshold $\delta_{s}$, the most recent $t$ number of deformation fields are deemed stable. TCI then proceeds to the subsequent convergence evaluation stage. Otherwise, another iteration is performed without the necessity to continue the subsequent stage.
@@
1:Input: $(l+1)$th deformation field ${\varphi}_{l+1}$, feature maps $F_{l}$, $M_{l}$, original images $I_{f}$, $I_{m}$, $W_{l}$, $k\leftarrow 1$
2:Output: $l$th deformation field ${\varphi}_{l}$
@@
5:  $M_{l}^{w}\leftarrow M_{l}\circ\varphi_{l}^{k-1}$#If $k=1\land l=4$, this step is skipped
6:  $R_{l}\leftarrow\left(|M_{l}^{w},F_{l}|\right)$ #If $k=1\land l=4$, $R_{4}\leftarrow\left(|M_{4},F_{4}|\right)$
7:  $O_{l}\leftarrow\text{{SEB}}(R_{l})$,  $H_{l}=\text{{SAB}}(|O_{l}|)$
8:  $O^{{}^{\prime}}_{l}\leftarrow O_{l}\ast H_{l}$ #AP output
9:  $F_{o}\leftarrow{Cl}_{2}(Res_{2}(Res_{1}(\gamma({Cl}_{1}(R_{l})))))$ #RP output
10:  $\overline{\varphi}_{l}^{k}={Cl}_{3}(|O^{{}^{\prime}}_{l},F_{o}|)$ #deformation estimation
11:  $\varphi_{l}^{k}\leftarrow{\varphi}_{l}^{k-1}\circ\overline{\varphi}_{l}^{k}+\overline{\varphi}_{l}^{k}$ #If $l=4\land k=1$, this step is skipped.
12:  $\tilde{\varphi}_{l}^{k}\leftarrow\text{upsample}(\varphi_{l}^{k},2^{l-1})$
13:  $s_{l}^{k}\leftarrow\text{sim}(I_{f},I_{m}\circ\tilde{\varphi}_{l}^{k})$ #see (14)
14:  append $s_{l}^{k}$ to $W_{l}$ #If $\text{len}(W_{l})=t+1$, remove $s_{l}^{k-t}$ from $W_{l}$
15:  if $\text{len}(W_{l})\geq t$ then#TCI invoked
16:   $\varepsilon_{l}\leftarrow\text{std}\left(W_{l}\right)$ #see (15)
17:   if $\varepsilon_{l}\leq\delta_{s}$ then #Stage 1: assess stability
18:    $\Delta s\leftarrow s_{l}^{k}-s_{l}^{k-1}$ #see (16)
19:    if $\Delta s\leq\delta_{c}$ then #Stage 2: assess convergence
20:     return ${\varphi}_{l}\leftarrow\text{upsample}(\varphi_{l}^{k},2)$
21:  $k\leftarrow k+1$
22:end while
23:return ${\varphi}_{l}\leftarrow\text{upsample}(\varphi_{l}^{k_{max}},2)$
@@
If $\Delta s$ falls below the predefined convergence threshold $\delta_{c}$, the deformation field at the $k$th iteration is deemed converged, the iteration terminates and the deformation field $\varphi_{l}$ is passed to the (l1)(l-1)th layer as input (see Figure 2). Otherwise, additional iterations continue at the $l$th layer. Following (Ma et al., 2024), TCI also leverages a predefined maximum number of iterations $k_{max}$ to avoid rare non-convergence situations. The pseudocode of the overall registration process is presented in Algorithm 1.
Notably, our TCI fundamentally differs from IIRP’s single‑stage method (Ma et al., 2024), which performs only one similarity check between two consecutive registrations and may prematurely terminate due to stochastic convergence. TCI adopts a two-stage evaluation with an additional stability check, enabling more informed stopping decisions while maintaining comparable inference time (see Table 4).
@@
Dice(%)↑ Time(s)↓ $|{J}_{\varphi}|_{\leq 0}$ Dice(%)↑ Time(s)↓ $|{J}_{\varphi}|_{\leq 0}$ Dice(%)↑ Time(s)↓ $|{J}_{\varphi}|_{\leq 0}$
U-Net VM (CVPR’18) 56.0 0.09 <<0.7% 65.1 0.02 <<0.3% 75.9 0.01 <<0.2% 319.4K
@@
NAP (BSPC’25) 66.8 0.30 <<0.2% 73.7 0.28 <<0.03% 80.4 0.22 <<0.07% 501.9K
[…]
3.3 Loss Functions
Being an unsupervised registration network, similar to the prior SOTA model (Ma et al., 2024), iPEAR employs NCC to evaluate the loss between the fixed image $I_{f}$ and the warped moving image $I_{w}^{m}$, which is defined as follows:
$\mathcal{L}_{ncc}(I_{f},I_{m}^{w})=-\sum\limits_{p\in\Omega}\frac{\sum\limits_{p_{i}}(I_{f}(p_{i})-\overline{I_{f}}(p))\cdot(I_{m}^{w}(p_{i})-\overline{I_{m}^{w}}(p))}{\sqrt{\sum\limits_{p_{i}}(I_{f}(p_{i})-\overline{I_{f}}(p))^{2}\cdot\sum\limits_{p_{i}}(I_{m}^{w}(p_{i})-\overline{I_{m}^{w}}(p))^{2}}},$
Moreover, to enforce the smoothness of the deformation field, we incorporate the $L_{2}$-norm $||\cdot||$ on the gradient of $\varphi$:
$\mathcal{L}_{smooth}(\varphi)=\sum\nolimits_{p\in\Omega}\lVert\nabla\varphi(p)\rVert^{2}.$ (18)
Therefore, the total loss in iPEAR is defined as follows:
$\mathcal{L}_{total}(I_{f},I_{m})=\mathcal{L}_{ncc}(I_{f},I_{m}^{w})+\lambda\mathcal{L}_{smooth}(\varphi),$ (19)
where $\lambda$ denotes the hyperparameter used to leverage relative importance in the overall loss function.
@@ 4 Experimental Results
To assess the performance of iPEAR, we conduct extensive experiments on three public MRI datasets, namely Mindboggle (Klein and Tourville, 2012), LPBA (Shattuck et al., 2008), and IXI (Avants et al., 2009), and one public abdomen CT dataset FLARE (Ma et al., 2022). While the SOTA models (Ma et al., 2024), (Cheng et al., 2025), and (Wang and Wang, 2025) were evaluated using two or three datasets, ours is more extensively evaluated. Appendix A describes these datasets, evaluation metrics, and experimental setups. The source code of iPEAR is available online anonymously [^1] https://anonymous.4open.science/r/iPEAR.
4.1 Comparison against Baseline Models
Figure 5: Visualization of registration results and corresponding deformation fields across different models on the Mindboggle dataset.
@@ 4.1 Comparison against Baseline Models
Model Dice(%)↑ Time(s)↓ $|{J}_{\varphi}|_{\leq 0}$
@@
NAP (BSPC’25) 74.9 0.16 <<2.0%
[…]
To benchmark the performance of iPEAR, we choose two U-Net models, namely VM (Balakrishnan et al., 2019) and TM (Chen et al., 2022), two cascaded models, namely RCN (Zhao et al., 2019a) and SDH (Zhou et al., 2023), and eight multi-scale pyramid models, namely RDN (Hu et al., 2022), PR++ (Kang et al., 2022), I2G (Liu et al., 2022), PIVIT (Ma et al., 2023), RDP (Wang et al., 2024), IIRP (Ma et al., 2024), SACB (Cheng et al., 2025), and NAP (Wang and Wang, 2025) as baseline models. For these baselines, we evaluate them using the same data preprocessing pipeline as iPEAR to ensure a fair comparison. To account for randomness, we evaluate each model three times with different random seeds on each dataset and report the averaged results. Unfortunately, we are not able to include certain recent baseline models (ReCorr (Li et al., 2025) and RDMR (Hu et al., 2025)) due to the incompleteness of the released source code. Table 1 presents the performance of different registration models on the Mindboggle, LPBA, and IXI datasets, measured using the Dice (Dice, 1945), Jacobian determinant ($|{J}_{\varphi}|_{\leq 0}$), and inference time (s) metrics. Appendix C presents HD95 (Huttenlocher et al., 1993), ASSD (Taha and Hanbury, 2015), and MSE metrics for these registration models across all three datasets. Table 2 presents the performance of different registration models (competing pyramid models in Table 1) on the FLARE dataset. The results shown in both tables demonstrate that iPEAR outperforms all the other registration models on the Dice metric across all four datasets, with particularly strong gains on Mindboggle, a well-known challenging benchmark with large spatial displacements. These gains are largely attributed to FARM, which mitigates the limitation of conventional residual convolutional decoders that are widely adopted by existing models (e.g., SACB). Such decoders treat feature channels uniformly and may miss subtle but critical deformation cues. Equipped with FARM, iPEAR produces more accurate and robust deformation estimates by emphasizing informative features and suppressing irrelevant ones. Furthermore, FARM is lightweight, enabling iPEAR to achieve superior performance with only a modest increase in parameter size (approximately 1.2×) over IIRP. Consequently, the inference time does not increase substantially. Unlike SACB, which estimates the deformation field only once per decoding layer, iPEAR employs TCI to adaptively determine the number of iterations at each layer, thereby yielding superior performance, achieving Dice improvements of 2.8%, 0.6%, 0.3%, and 2.6% on the four datasets, respectively. Notably, RDP and IIRP also employ iterative strategies. However, RDP adopts a fixed number of iterations and IIRP relies on a single-stage stopping criterion evaluation, both of which lead to inferior performance. In contrast, our iPEAR employs TCI with a dual-stage strategy at each decoding layer, thereby achieving elevated performance while incurring only a marginal increase in inference time over IIRP.
Figure 5 and Appendix D visualize the registration results on the Mindboggle and LPBA datasets, respectively. On Mindboggle, performance differences among models are more noticeable in the two structural regions of the brain images, highlighted using the red and yellow boxes. These visualizations further demonstrate that the moving images, after being warped by the deformation fields produced by iPEAR, align more closely with fixed images. Additionally, we perform quantitative analyses of specific regions in the left and right hemispheres to further illustrate variations in anatomical structure registration performance across different models. The results, as presented in Appendix E, further demonstrate the superior performance of iPEAR.
4.2 Generalizability of FARM and TCI
Table 3: The generalizability of FARM on the IXI dataset
Model Dice(%) ↑ MSE($10^{-3}$) ↓
I2G (+ FARM) 79.3 (79.9) 0.40 (0.37)
PR++ (+ FARM) 79.3 (79.4) 0.36 (0.32)
IIRP (+ FARM) 80.4 (80.5) 0.24 (0.20)